OpenStudy (anonymous):

Fan and Medal ! Explain how quadratic trinomial can be factored by grouping. Include examples in your explanation.

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

4x^3 + 4x^2 + 4x + 4 = 0 This is a 4 term polynomial 4x^2(x + 1) + 4(x + 1) = 0 (4x^2 + 4)(x + 1) = 0 That's how it's factored. i don't know about the quadratic trinomial.

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

Now I just need the quadratic trinomial

OpenStudy (anonymous):

Hows this ? a three term trinomial can be made into a four term polynomial. 4x^2+7x+3=4x^2+4x+3x+3

OpenStudy (johnweldon1993):

Oh...well a quadratic trinomial is just a polynomial with a highest degree of 2 so \[\large x^2 - x + 2\] Would be an example... Factoring this would just be \[\large x^2 - x + 2 = 0\] \[\large (x -2)(x - 1) = 0\]

OpenStudy (johnweldon1993):

And above what you posted last...is another way to factor it as well...now you would be able to factor by grouping that way....but the method you have up at the top is correct as well

OpenStudy (anonymous):

ok cool thanks, can you help me with something else ?